Based upon your estimated purity of the E3 fraction in the question above, and upon knowing the amount of protein present in your fraction, estimate the total yield (in ug) of rGFP obtained using this protocol. You must show/explain how you calculated your answer.
Answer #1

Methodology of Purification of RGFP Biology Essay

1. Expression in Bacteria

For the expression of bacteria, 2 bacterial cultures were used. One denoted as V, BL21,DE3, pLysS, pRSETA. The other denoted as G, BL21, DE3, pLysS, pRSETA-GFPuv. These were inoculated with LB (100ug/ml of ampicillin. 25ug/ml of cam).

2. Preparation of crude extract should be done for bacterial culture.

3. after that Ni 2+ Agarose Affinity Chromatography should be done.

4. Later bradford assay shoulbe done to mesure the total protein at 595 nanometer wave length.

5. SDS PAGE should be done for analysis of rGFP fraction.

6. Finally western blot analysis should be done for analysis of desired protein fraction.

Calculation of Bradford analysis:

Graph Absorbance vs concentration, and obtains the equation of the line (y = mx + b), with r2, as close to 1 as possible.

Another option is to subtract the value of "zero" protein, then the straight line passes through zero and the equation simplifies: y = mx.

Where y is the absorbance, m is the slope and x is the concentration.

Thus, x = y / m.

The absorbance of the sample is divided by the value of the slope and the concentration is obtained.

If the sample was diluted, the value is adjusted by multiplying by the dilution factor.

Example:

Absorbance = 0. 36
Sample was diluted 10 times
The absorbance of the "zero" is 0.053
Adjusted absorbance: 0. 360-0.053 = 0.307

The standard curve equation: y = 0.012x

x = y / m, therefore x = 0.307 / 0.012
x = 25.58 ug / ml

But sample diluted 10 times:

x = 25.58 x 10 = 255.8 ug prot / ml

The results will be able to express rGFP in the E. coli bacterial strain. This is evident because of the Bradford Assay, SDS-PAGE gel, and Western Blot analysis of our Ni 2+ Agarose chromatography purification step. The Bradford Assay let us determine the total amount of protein in each of the fractions. From there the SDS-PAGE gel will give us insight as to just how much rGFP will be in the fractions, more specifically the E3 fraction, by determining how pure bands are based on molecular weight. The Western Blot continued this type of analysis based on molecular weight, but it also use the binding of antibodies onto the Xpress Epitope on the rGFP strain.